# Implicit Bias of the JKO Scheme

### v1

@@ 1. Introduction
$W_{g}^{2}(\mu,\nu):=$ $\inf_{\gamma\in\Gamma(\mu,\nu)}\iint_{M\times M}d_{g}^{2}(x,y)\gamma(dx,dy),$
$d_{g}(x,y)$ denotes the Riemannian distance between $x,y\in M$ and the infimum is over all couplings between $\mu,\nu$, the set of all $\gamma\in P_{ac}(M\times M)$ with marginals $\int_{M}\gamma(x,dy)=\mu(x)$ and $\int_{M}\gamma(dx,y)=\nu(y)$. A key result in [AGS2008] is that the proximal point scheme (4), which is often called the JKO scheme, is a first order integrator of the Wasserstein gradient flow.
Theorem 1 ([AGS2008] Theorem 4.0.4 v).
Suppose $J$ is $\lambda$-convex and smooth. Then for any $T>0$
$\sup_{t\in[0,T]}W_{2}\left(\rho_{t},\rho_{\lfloor t/\eta\rfloor}^{\mathrm{JKO},\eta}\right)=O_{T}(\eta).$
@@ 1.1. Main Result: Implicit Bias of JKO
    $J$ is proper, l.s.c., and $\lambda$-geodesically convex in $(P_{2},W_{2})$ for $\lambda>0$
@@
    $\partial_{t}\rho^{\mathrm{GF}}_{t}+\nabla\cdot(-\nabla\psi(\rho_{t}^{\mathrm{GF}})\rho_{t}^{\mathrm{GF}})=0,\qquad\psi=\frac{\partial J}{\partial\rho}$
    admits classical solution $(\rho^{\mathrm{GF}}_{t})\in C_{t,x}^{2}$ on $t/\eta\in[n,(n+1)]$. Further, suppose $\rho_{t}$ satisfies fast decay to zero at infinity so integration by parts $\int\varphi\partial_{t}\rho_{t}=-\int\varphi\,\mathrm{div}(\rho_{t}v_{t})=\int\langle\nabla\varphi,v_{t}\rangle\rho_{t}$ is valid on $M$.
@@ 1.2. Illustrative Examples
We further remark on other implicit biases of interest in Section B.1, including interaction energies, the Porous-Medium equation, and $K$-Wasserstein barycenter.
@@ 1.5. Implicit Bias of JKO on the Free-Energy Functional and Langevin Dynamics
(Remark) The Bohm Quantum potential $Q[\rho]=-\frac{\hbar^{2}}{2m}\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}$ remarkably arises as a scaling of the first variation of the Fisher functional, with applications in Quantum-Drift-Diffusion (QDD) [Gianazza2008]. This offers insight into the mysterious term (24). $Q[\rho]$ is a global, non-local penalty on curvature of the density which differentiates between classical and quantum mechanics. Viewed in this sense, this term of the implicit bias exhibits a global, non-local regularization of the curvature of the entire density with ’quantum-mechanical’ behavior.
@@ 2. Strategy of Proof of Theorem 2
${\bm{j}}(\rho)=\nabla_{g}\frac{\delta\left|\partial J\right|^{2}}{\delta\rho}(\rho),\qquad\left|\partial J\right|^{2}=\frac{1}{4}\int_{M}\left\lVert\nabla_{g}\frac{\partial J}{\partial\rho}\right\rVert^{2}\rho(dx).$
@@ Appendix A Other Background
Wasserstein Gradient Flows and the JKO Scheme. The Jordan-Kinderlehrer-Otto Scheme [Jordan1998], and its continuous-time limit the Wasserstein gradient flow [Jordan1998, AGS2008], are a general approach for optimization problems over distributions and PDEs. Their universality is reflected in the number of problems they recover – from stochastic differential equations following Langevin dynamics [Jordan1998], the heat equation [Jordan1998], filtering problems [Halder2017], Wasserstein geodesics and optimal transport [peyre2019computational], and so on, this framework yields one of the most canonical means of minimizing an energy functional defined over distributions. In Wasserstein Gradient Flow we are given a functional $J:{P}_{ac}(\mathbb{R}^{d})\to\mathbb{R}$, the measure $(\rho_{t})_{t\,\geq\,0}$ is the Wasserstein gradient flow of $J$ if it solves the continuity equation with respect to the gradient of the first-variation (Wasserstein gradient) of $J$:
$\partial_{t}\rho_{t}=+\nabla\cdot\left(\rho_{t}\nabla\frac{\delta J}{\delta\rho}\right)$ (41)
The Jordan-Kinderlehrer-Otto Scheme [Jordan1998] is akin to an implicit Euler discretization of the Wasserstein gradient flow, defined by the optimization problem over square-integrable, absolutely continuous distributions $\rho\in{P}_{ac}(\mathbb{R}^{d})$
$\rho^{(k+1)}\leftarrow\operatorname*{arg\,min}_{\rho\in{P}_{ac}(\mathbb{R}^{d})}\,\,J(\rho)+\frac{1}{2\eta}W_{2}^{2}(\rho^{(k)},\rho)$ (42)
In particular, as any backward Euler discretization scheme, (4) corresponds to a first order integrator of the Wasserstein gradient flow.
Theorem 4 ([Jordan1998, AGS2008]).
Suppose $J$ is $\lambda$-convex and smooth. Then for any $T>0$
$\sup_{t\in[0,T]}W_{2}\left(\rho_{t},\rho_{\lfloor t/\eta\rfloor}^{JKO,J,\eta}\right)=O_{T}(\eta).$
Notably, (41) is often tractable to solve with particle schemes for many classes of energy functionals. However, (4) is often impossible to solve directly, with the exception of special cases such as Linear SDEs and Gaussian filtering [Halder2017], owing to the difficulty of the infinite-dimensional optimization problem.
@@ Appendix B Implicit Bias of the JKO Scheme
    $J$ is proper, l.s.c., and $\lambda$-geodesically convex in $(\mathbb{P}_{2},W_{2})$.
@@
    admits classical solution $(\rho^{\mathrm{GF}}_{t})\in C_{t,x}^{2}$ on $t/\eta\in[n,(n+1)]$. Further, suppose $\rho_{t}$ satisfies fast decay to zero at infinity so integration by parts is valid.
@@
We switch signage conventions on $\bm{j}_{t}$ in  53 below, to be consistent with the the forward-direction of [smith2021on] in explicit Euler.
@@
We briefly show  4. Let us suppose, as standard, a compact support $\Omega$ with vanishing boundary conditions. Further, suppose we take compact variations of the form $\rho_{(s)}=\rho+s\dot{\phi}$. One first simply takes the $s$-derivative
@@
Suppose $\eta>0$ is a step-size and $\beta>0$ is an inverse-temperature. Then, minimizing the energy 65 over $\rho\in\mathrm{BW}(\mathbb{R}^{d})$ with (4) coincides with the following Bures-Wasserstein gradient-flow up to $O_{T}(\eta^{2})$
@@ Appendix D Implicit Regularization of Riemannian Gradient Descent
While Riemannian gradient flow is guaranteed to remain on $M$, unlike in the case of Euclidean gradient descent, the discretization of 74 given by $p^{(k+1)}=p^{(k)}-\eta\nabla_{\!g}J$ generally fails to remain. To formalize a discretization which ensures feasibility, for any $p,q\in M$ denote the Riemannian geodesic distance between the points by $d_{M}(p,q)$ by the infimal length of a curve $\gamma\in C^{\infty}[0,1]$ parametrized by arc length which satisfies $\gamma(0)=p$ and $\gamma(1)=q$.
@@
In Riemannian gradient descent one minimizes 73 by choosing a step-size $\eta\in\mathbb{R}_{+}$ and solving the following proximal-point implicit step for iterations $k\in[T]$
@@
Which ensures that the sequence of iterates $(p^{(t)})_{t=1}^{T}\subset M$. An equivalent explicit form for this step is given by the exponential map. Let $(p,v)\in TM$ denote a point in the tangent bundle. The exponential map $\mathrm{exp}_{p}(v):TM\to M$ is defined by $\mathrm{exp}_{p}(v)=\gamma_{v}(1)$ where $\gamma_{v}:[0,1]\to M$ is the unique geodesic p.a.l. satisfying $\gamma_{v}(0)=p$ and $\gamma_{v}^{\prime}(0)=v$. Given this definition, the equivalent explicit step for 76 is given by the update

### v2

@@ 1. Introduction
$W_{g}^{2}(\mu,\nu):=$ $\inf_{\gamma\in\Gamma(\mu,\nu)}\int_{M\times M}d_{g}^{2}(x,y)\gamma(dx,dy),$
$d_{g}(x,y)$ denotes the Riemannian distance between a pair of points $x,y\in M$ and the infimum is over all couplings between $\mu$ and $\nu$, the set of all $\gamma\in P_{ac}(M\times M)$ with marginals $\int_{M}\gamma(x,dy)=\mu(x)$ and $\int_{M}\gamma(dx,y)=\nu(y)$ denoted by $\Gamma(\mu,\nu)$.
Definition 1 (Order-$k$ Integrator for the JKO Scheme.).
Suppose $J:P_{ac}\to\mathbb{R}$ is an energy functional for which the JKO Scheme is well-posed and stable with piecewise interpolation $\rho_{\lfloor t/\eta\rfloor}^{\mathrm{JKO}}$. A curve $(\rho^{\prime}_{t})_{t\geq 0}\subset P_{ac}$ is called an order-$k$ integrator for the JKO scheme if for any $T>0$
$\sup_{t\in[0,T]}W_{2}(\rho^{\prime}_{t},\rho_{\lfloor t/\eta\rfloor}^{\mathrm{JKO},\eta})=O_{T}(\eta^{k}).$
A key result in [AGS2008] is that the proximal point scheme (4), which is often called the JKO scheme, is a first order integrator of the Wasserstein gradient flow of $J$.
Theorem 1 ([AGS2008] Theorem 4.0.4, 11.2.1).
Suppose $J$ is proper, l.s.c., coercive, and $\lambda$-geodesically convex along generalized geodesics. Then the JKO Scheme is well-posed and stable, with piecewise interpolation $\rho_{\lfloor t/\eta\rfloor}^{\mathrm{JKO}}$. Moreover, the Wasserstein gradient flow $(\rho_{t})_{t\geq 0}$ defined by
$\partial_{t}\rho_{t}=\mathrm{div}\left(\rho_{t}\nabla\frac{\delta J}{\delta\rho}(\rho_{t})\right),$
exists, is unique, and satisfies for any $T>0$
$\sup_{t\in[0,T]}W_{2}(\rho_{t},\rho_{\lfloor t/\eta\rfloor}^{\mathrm{JKO},\eta})=O_{T}(\eta).$
Theorem 1 is usually interpreted in the forward direction: the JKO Scheme approximates the Wasserstein gradient flow with $O_{T}(\eta)$ accuracy. In reverse, in the sense of Definition 1, the Wasserstein gradient flow $(\rho_{t})_{t\geq 0}$ is itself an order-1 integrator for the JKO Scheme.
In this work, we take this first-order consistency estimate as a black-box: we assume that the energy functional $J$ satisfies the hypotheses guaranteeing the existence of the gradient flow, the well-posedness of the JKO Scheme, and the stability and energy-dissipation estimates of [AGS2008] (e.g. the assumptions of Theorem 11.2.1). Under these conditions, the JKO Scheme provides an $O_{T}(\eta)$ approximation to the continuous flow. Our main result builds on this first-order relationship by identifying a second-order modified flow matching the JKO iterates to $O_{T}(\eta^{2})$.
@@ 1.1. Main Result: Implicit Bias of JKO
    $J$ is proper and l.s.c. in $(P_{2},W_{2})$.
@@
    $\partial_{t}\rho^{\mathrm{GF}}_{t}+\nabla\cdot(-\nabla\psi(\rho_{t}^{\mathrm{GF}})\rho_{t}^{\mathrm{GF}})=0,\qquad\psi=\frac{\delta J}{\delta\rho}$
    admits classical solution $(\rho^{\mathrm{GF}}_{t})\in C_{t,x}^{2}$ on $t/\eta\in[n,(n+1)]$ which is $O_{T}(\eta)$ from $\rho_{k}^{\mathrm{JKO},\eta}$ (Theorem 1). Further, suppose $\rho_{t}$ satisfies fast decay to zero at infinity so integration by parts $\int\varphi\partial_{t}\rho_{t}=-\int\varphi\,\mathrm{div}(\rho_{t}v_{t})=\int\langle\nabla\varphi,v_{t}\rangle\rho_{t}$ is valid on $M$.
@@
Theorem 2, is a local, second-order result. It reverses the usual direction of the analysis: instead of showing the JKO Scheme approximates a given flow, it identifies a modified energy $J^{\eta}$ whose Wasserstein gradient flow coincides with the discrete JKO updates to $O_{T}(\eta^{2})$. This construction only requires local smoothness and well-posedness assumptions on $J$ and on the continuous and discrete flows; it does not explicitly require $\lambda$-geodesic convexity.
Although $\lambda$-convexity (with properness, l.s.c., and coercivity) is not needed for this second-order analysis, it is a standard sufficient condition which [AGS2008] guarantees existence, uniqueness, stability, and first-order consistency of both the Wasserstein gradient flow and JKO scheme (Theorem 1).
@@ 1.2. Illustrative Examples
We further remark on other implicit biases of interest in Section B.1, including interaction energies, the Porous-Medium equation, and $K$-Wasserstein barycenter. Notably, with the implicit Euler signage the correction $\frac{\eta}{4}|\partial J(\rho)|^{2}$ is concave in $W_{2}$. A natural question, then, is to determine for which $J$ and ranges of $\eta$ the modified Wasserstein gradient flow generated by $J^{\eta}$ remains well-posed. While we compute the induced regularization of the JKO Scheme for several examples, we do not attempt here to prove full well-posedness for each case.
@@ 1.5. Implicit Bias of JKO on the Free-Energy Functional and Langevin Dynamics
(Remark) The Bohm Quantum potential $Q[\rho]=-\frac{\hbar^{2}}{2m}\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}$ remarkably arises as a scaling of the first variation of the Fisher functional, with applications in Quantum-Drift-Diffusion (QDD) [Gianazza2008]. This offers insight into the mysterious term (24). $Q[\rho]$ is a global, non-local penalty on curvature of the density which differentiates between classical and quantum mechanics. Viewed in this sense, this term of the implicit bias exhibits a global, non-local regularization of the curvature of the entire density with “quantum-mechanical” behavior.
@@ 2. Strategy of Proof of Theorem 2
${\bm{j}}(\rho)=\nabla_{g}\frac{\delta\left|\partial J\right|^{2}}{\delta\rho}(\rho),\qquad\left|\partial J\right|^{2}=\frac{1}{4}\int_{M}\left\lVert\nabla_{g}\frac{\delta J}{\delta\rho}\right\rVert^{2}\rho(dx).$
@@ Appendix B Implicit Bias of the JKO Scheme
    $J$ is proper and l.s.c. in $(\mathbb{P}_{2},W_{2})$.
@@
    admits classical solution $(\rho^{\mathrm{GF}}_{t})\in C_{t,x}^{2}$ on $t/\eta\in[n,(n+1)]$ which is $O_{T}(\eta)$ from $\rho_{k}^{\mathrm{JKO},\eta}$ (Theorem 1). Further, suppose $\rho_{t}$ satisfies fast decay to zero at infinity so integration by parts is valid.
@@
Remark 2.
We note that requiring $J$ to be $\lambda$-geodesically convex (with sufficient coercivity) in addition to (A1) is not necessary for this local result to hold. However, if one does assume $\lambda$-convexity, the theory of metric gradient flows [AGS2008] guarantees $J$ has a unique, well-behaved Wasserstein gradient flow, and that the JKO scheme is well-posed, stable, and converges to this flow. These properties ensure the regularity assumptions (A2) and (A4).
@@
We switch signage conventions on $\bm{j}_{t}$ in  51 below, to be consistent with the the forward-direction of [smith2021on] in explicit Euler.
@@
We briefly show  5. Let us suppose, as standard, a compact support $\Omega$ with vanishing boundary conditions. Further, suppose we take compact variations of the form $\rho_{(s)}=\rho+s\dot{\phi}$. One first simply takes the $s$-derivative
@@
Suppose $\eta>0$ is a step-size and $\beta>0$ is an inverse-temperature. Then, minimizing the energy 63 over $\rho\in\mathrm{BW}(\mathbb{R}^{d})$ with (4) coincides with the following Bures-Wasserstein gradient-flow up to $O_{T}(\eta^{2})$
@@ Appendix D Implicit Regularization of Riemannian Gradient Descent
While Riemannian gradient flow is guaranteed to remain on $M$, unlike in the case of Euclidean gradient descent, the discretization of 72 given by $p^{(k+1)}=p^{(k)}-\eta\nabla_{\!g}J$ generally fails to remain. To formalize a discretization which ensures feasibility, for any $p,q\in M$ denote the Riemannian geodesic distance between the points by $d_{M}(p,q)$ by the infimal length of a curve $\gamma\in C^{\infty}[0,1]$ parametrized by arc length which satisfies $\gamma(0)=p$ and $\gamma(1)=q$.
@@
In Riemannian gradient descent one minimizes 71 by choosing a step-size $\eta\in\mathbb{R}_{+}$ and solving the following proximal-point implicit step for iterations $k\in[T]$
@@
Which ensures that the sequence of iterates $(p^{(t)})_{t=1}^{T}\subset M$. An equivalent explicit form for this step is given by the exponential map. Let $(p,v)\in TM$ denote a point in the tangent bundle. The exponential map $\mathrm{exp}_{p}(v):TM\to M$ is defined by $\mathrm{exp}_{p}(v)=\gamma_{v}(1)$ where $\gamma_{v}:[0,1]\to M$ is the unique geodesic p.a.l. satisfying $\gamma_{v}(0)=p$ and $\gamma_{v}^{\prime}(0)=v$. Given this definition, the equivalent explicit step for 74 is given by the update